OpenStudy (yeval76):

***FAN AND MEDAL*** You decide to put $150 in a savings account to save for a $3,000 down payment on a new car. If the account has an interest rate of 2.5% per year and is compounded monthly, how long does it take you to earn $3,000 without depositing any additional funds? 10.1101 years 119.954 years 121.321 years 134.34 years

OpenStudy (yeval76):

@m_akram

OpenStudy (yeval76):

@dan815

OpenStudy (yeval76):

@StudyGurl14 help please

OpenStudy (studygurl14):

\(\large A=P(1+\Large\frac{r}{n})^{nt}\)

OpenStudy (studygurl14):

A = total money at the end of period P = principal, or beginning amount of money r = rate n = amount of compounds per year (in this case, 12) t = number of years Plug in 3000 for A, and all the rest of the known values

OpenStudy (studygurl14):

and solve for t

OpenStudy (anonymous):

which means that the t is going to be 119.954

OpenStudy (yeval76):

ok so it would be \[3,000 = 150(1 + \frac{ .025 }{ 12 })^{12t}\]

OpenStudy (studygurl14):

yes

OpenStudy (studygurl14):

simplify it

OpenStudy (studygurl14):

and solve for t

OpenStudy (yeval76):

so then t would equal 12.10732232 @StudyGurl14

OpenStudy (yeval76):

@StudyGurl14

OpenStudy (yeval76):

but thats not in my answers

OpenStudy (studygurl14):

how about, instead of trying to solve for t, you just plug in the answer choices into the equation until you get a value of A that is at least 3000

OpenStudy (yeval76):

ok so its a

OpenStudy (yeval76):

*A 10.1101 @StudyGurl14

OpenStudy (retireed):

the answer is 119.9540698 years B

OpenStudy (studygurl14):

No, a is not right

OpenStudy (studygurl14):

It is way off

OpenStudy (yeval76):

THANK YOU!!! I didn't really get it at first...

OpenStudy (studygurl14):

Yes, B is correct. Did you try plugging in to see which answer choice got approx 3000?

OpenStudy (yeval76):

Yes but both told me 3000

OpenStudy (retireed):

Who's way off ? It's B